Shift report 20011016 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 16 2001 - 16:01:20 EDT

  • Next message: Zbigniew Majka: "Odp: DC"

    ------------------------------------------------------------
    Date: Tue Oct 16 16:01:20 2001, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#10 pages 17 - 17
    
    Shift Summary:
    ------------------------------------------------------------
    1.  Beam for 1/2 hours. Machine had "cryo system problems". 
        (20 min. in the morning, 10 min. in the afternoon.)
        At 15:51,  beam was stored and cogged.
    
    2.  Not much...  
        Data Run :  5370 (70k trigger 1)
                    5374 (26k trigger 1)
                    5378 (start 15:59, and now running) 
    
        Pedestal & DVM Run : 5375, 5376, 5377   
    
    3.  No new detector problems.
    
        The problems in previous shift report :
         - ZDC Right Module 1 Low : FV fixed.
           -> switched the "capacitance coupling cable"
         - BB ADC Left Module 37 : EJK turned on HV.
    
        FV added trigger 5 plots to SummaryPlots.C
    
        
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 16 2001 - 16:02:09 EDT